• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Yi Qiang

    Log entry time 20:42:33 on April 22, 2008

    Entry number 233820

    This entry is a followup to: 233815

    keyword=A detailed description of target after the power trip

    After the power trip, the High Power Heater of loop1 went to 600W (its maximum value) and the target temperature and target vacuum increased repidly. The fan speed in Loop1 also went wrong. So we turned off the HPH, moved the cryo target away from beam and changed the fan speed back to 30 Hz. After a while, as the target temperature and pressure got stabilized, we tried to turn on the HPH again and now it looked normal. Since nothing is abnormal now, after we finish the Al dummy test, we will continue to take production data on LH2.